# A Behaviour and Disease Model of Testing and Isolation

### v1

@@ 4.1.3 Effect of behavioural uptake rates on the dynamics
Spontaneous uptake of behaviour ($\omega_{3}$) has a similar effect as social pressure (Figure 4, (c), (f)). Increasing spontaneous uptake in the model reduces the true epidemic size while increasing the peak observed number of cases. Removal of spontaneous uptake ensures that the epidemic is not observed in this model. Due to the baseline parameter values (Table 1), without a small influence of spontaneous behavioural uptake there is not enough awareness of testing in the population for testing to “take off”, leading to a severe lack of testing and reporting of the epidemic. In this setting, we would observe a silent epidemic where many individuals become unwell but do not seek testing and isolation. Increasing spontaneous uptake sufficiently has the same effect in controlling the epidemic as social pressure. Like social pressure, this control of the epidemic is driven by the speed of behaviour uptake and the increase in $B(0)$ (Figure 4 (i)).
@@
The effect of behavioural parameters on the transient dynamics are similar in an endemic regime, albeit with trajectories converging to steady states (Figure S3).
@@ 4.2 Endemic regime with waning immunity
Here we detail our findings for an endemic regime with $\nu>0$ as detailed in Table 1. We report on the impact of testing and isolation effects on endemic steady states, the impact of behaviour and disease contagiousness on endemic states, and the stability of steady states.
@@ 4.2.1 Testing and isolation effects on endemic steady states
Figure 5: The relationship between isolation & test effectiveness and endemic states. In each panel we display outputs for different combinations of isolation effectiveness ($1-q_{T}$, y-axis) and test effectiveness ($p_{T}$, x-axis). In all simulations we used a fixed behaviour-free reproduction number of $\mathcal{R}_{0}^{D}=3.28$. (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population ($I$); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) Basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number.
@@ 4.2.2 Behaviour and disease contagiousness on endemic states
Figure 6: The effect of the social reproduction number and the behaviour-free reproduction number on endemic states. In each panel we display outputs for different combinations of the social reproduction number ($\mathcal{R}_{0}^{B}$, $y$-axis) and the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$, $x$-axis). The measures displayed are: (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population ($I$); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) The basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameters in Table 1.
@@ 4.2.3 Stability of steady states
There are four feasible steady state combinations for this BaD model for testing and isolation: no behaviour or disease (denoted $E_{00}$); no behaviour, disease endemic (denoted $E_{0D}$); behaviour endemic, no disease (denoted $E_{B0}$), and; behaviour and disease endemic (denoted $E_{BD}$). The stable regions of the steady states agree with the regions previously identified in [17] when $\omega_{1},\omega_{2},\omega_{3}>0$ (Figure 7 (a)) and when $\omega_{1}>0,\omega_{2}=\omega_{3}=0$ (Figure 7 (c)). In contrast, when $\omega_{3}=0$ and $\omega_{1},\omega_{2}>0$ we observe a non-linear boundary where the stable steady state switches from $E_{0D}$ to $E_{BD}$ (Figure 7 (b)). This boundary is driven by a non-linear relationship between social spread of behaviour ($\mathcal{R}_{0}^{B}$) and the perception of illness threat ($\omega_{2}$). This suggests that there are parameter combinations where an infection can be endemic in a population without being detectable in the sense that negligibly few tests are being conducted. Since those who test positive ($T$) are explicitly a subset of those willing to test and isolate ($B$), the threshold observed in Figure 7 (b) shows there is a need for enough people actively to seek testing and isolate if positive before the feedback mechanism of the perception of illness threat can take effect.
@@ 5.2 Study limitations and directions for further work
We regard the model we have presented as an entry point for quantitative study of the feedbacks between testing behaviour and infection dynamics. As with any model we acknowledge that it is, by necessity, a simplified representation of reality. It is important that we consider the modelling assumptions made and their potential limitations. Here we expand on the implications of our behavioural parameter assumptions, testing assumptions and reduction of model dimensionality via assuming the pathogen to have a single strain and there being a sole testing intervention deployed.
A core aspect of our presented model is parsimony. We therefore had fixed values for the behaviour effect parameters in the testing behaviour uptake rate ($\omega_{1},\omega_{2},\omega_{3}$) and testing behaviour abandonment rate ($\alpha_{1},\alpha_{2}$). Relaxing this assumption would permit temporal variation in these behavioural factors. One scenario of interest would be to have a decay rate on the perception of illness threat ($\omega_{2}$), corresponding to growing normality through time amongst the population for cases occurring and infection circulating in the community. Alternatively, we could vary the epidemiological metric to which the perception of illness threat is tied. We associated $\omega_{2}$ with those who tested positive. However, as there may be delays in reporting the count of positive tests to the wider population, different lags to this measure could be applied. Further, the case severity that the population responds to may alter during an outbreak (i.e. at different times the number of cases, hospitalisations and/or deaths may be attributed different importance); consideration of these behavioural feedbacks would require refinements to our model to include severe case outcomes (such as hospitalisations and deaths).
@@ Acknowledgements
EMH is funded by The Pandemic Institute, formed of seven founding partners: The University of Liverpool, Liverpool School of Tropical Medicine, Liverpool John Moores University, Liverpool City Council, Liverpool City Region Combined Authority, Liverpool University Hospital Foundation Trust, and Knowledge Quarter Liverpool (EMH is based at The University of Liverpool). The views expressed are those of the author(s) and not necessarily those of The Pandemic Institute. TH is supported by the Wellcome Trust (Grant number 227438/Z/23/Z). MGR is supported by the Marsden Fund under contract 20-MAU-50. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme “Modelling and inference for pandemic preparedness”, where work on this paper was initiated. This work was supported by EPSRC grant EP/Z000580/1. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 4.1.3 Effect of behavioural uptake rates on the dynamics
Spontaneous uptake of behaviour ($\omega_{3}$) has a similar effect as social pressure (Figure 4, (c), (f)). Increasing spontaneous uptake in the model reduces the true epidemic size while increasing the peak observed number of cases. Removal of spontaneous uptake ensures that the epidemic is not observed in this model. Due to the baseline parameter values (Table 1), without a small influence of spontaneous behavioural uptake there is not enough awareness of testing in the population for testing to “take off”, leading to a severe lack of testing and reporting of the epidemic. In this setting, we would observe a silent epidemic where many individuals become unwell but do not seek testing and isolation. Increasing spontaneous uptake sufficiently has the same effect in controlling the epidemic as social pressure. Like social pressure, this control of the epidemic is driven by the speed of behavioural uptake and the increase in $B(0)$ (Figure 4 (i)).
@@
The effect of behavioural parameters on the transient dynamics are similar in an endemic regime, albeit with trajectories converging to steady states (Supplementary Figure 11).
@@ 4.2 Endemic regime with waning immunity
Here we detail our findings for an endemic regime with $\nu>0$ as detailed in Table˜1. We report on the impact of testing and isolation effects on endemic steady states, the impact of behaviour and disease contagiousness on endemic states, and the stability of steady states.
@@ 4.2.1 Testing and isolation effects on endemic steady states
Figure 5: The effect of isolation and test effectiveness on endemic states. In each panel we display outputs for different combinations of isolation effectiveness ($1-q_{T}$, y-axis) and test effectiveness ($p_{T}$, x-axis). In all simulations we used a fixed behaviour-free reproduction number of $\mathcal{R}_{0}^{D}=3.28$. (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population (II); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) Basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number.
@@ 4.2.2 Behaviour and disease contagiousness on endemic states
Figure 6: The effect of the social reproduction number and the behaviour-free reproduction number on endemic states. In each panel we display outputs for different combinations of the social reproduction number ($\mathcal{R}_{0}^{B}$, $y$-axis) and the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$, $x$-axis). The measures displayed are: (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population (II); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) The basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameters in Table 1.
@@ 4.2.3 Stability of steady states
There are four feasible steady state combinations for this BaD model for testing and isolation: no behaviour or disease (denoted $E_{00}$); no behaviour, disease endemic (denoted $E_{0D}$); behaviour endemic, no disease (denoted $E_{B0}$), and; behaviour and disease endemic (denoted $E_{BD}$). The stable regions of the steady states agree with the regions previously identified by Ryan et al. [19] when $\omega_{1},\omega_{2},\omega_{3}>0$ (Figure 7 (a)) and when $\omega_{1}>0,\omega_{2}=\omega_{3}=0$ (Figure 7 (c)). In contrast, when $\omega_{3}=0$ and $\omega_{1},\omega_{2}>0$ we observe a non-linear boundary where the stable steady state switches from $E_{0D}$ to $E_{BD}$ (Figure 7 (b)). This boundary is driven by a non-linear relationship between social spread of behaviour ($\mathcal{R}_{0}^{B}$) and the perception of illness threat ($\omega_{2}$). This suggests that there are parameter combinations where an infection can be endemic in a population without being detectable in the sense that negligibly few tests are being conducted. Since those who test positive ($T$) are explicitly a subset of those willing to test and isolate ($B$), the threshold observed in Figure 7 (b) shows there is a need for enough people actively to seek testing and isolate if positive before the feedback mechanism of the perception of illness threat can take effect.
@@ 5.1 Findings in context
Compared to the original BaD model [19], we observed similarities in the thresholds and steady states. Despite the obvious complexity induced by a more complicated disease structure, we found here that the reproduction number can also be decomposed into four components describing the relative contribution for the mixing behavioural groups. Further, Theorem 3.2 shows a similar complexity arising in the steady states, with analytic expressions in terms of model states found numerically. However, there are interesting differences in the dynamics and steady state properties. In particular, the role of the perception of illness threat is different. Here, we have modelled the perception of illness threat directly on the prevalence of positive tests in the population, as opposed to a less well-defined prevalence of cases in the original model. This has caused an interesting threshold in the absence of spontaneous behavioural uptake (Figure 7 (b)). If testing behaviour is not socially contagious enough, we find that there exists a disease endemic state with no testing, which is not possible in the original BaD model. The influence of perception of illness threat also creates the possibility of epidemics with false elimination (the hill structure in Figure 6 (c)).
@@ 5.2 Study limitations and directions for further work
We regard the model we have presented as an entry point for quantitative study of the feedbacks between symptomatic testing behaviour and infection dynamics. As with any model we acknowledge that it is, by necessity, a simplified representation of reality. It is important that we consider the modelling assumptions made and their potential limitations. Here we expand on the implications of our behavioural parameter assumptions, testing assumptions and reduction of model dimensionality via assuming the pathogen to have a single strain and there being a sole testing intervention deployed.
A core aspect of our presented model is parsimony. We therefore had fixed values for the behaviour effect parameters in the testing behavioural uptake rate ($\omega_{1},\omega_{2},\omega_{3}$) and testing behaviour abandonment rate ($\alpha_{1},\alpha_{2}$). Relaxing this assumption would permit temporal variation in these behavioural factors. One scenario of interest would be to have a decay rate on the perception of illness threat ($\omega_{2}$), corresponding to growing normality through time amongst the population for cases occurring and infection circulating in the community. Alternatively, we could vary the epidemiological metric to which the perception of illness threat is tied. We associated $\omega_{2}$ with those who tested positive. However, as there may be delays in reporting the count of positive tests to the wider population, different lags to this measure could be applied. Further, the case severity that the population responds to may alter during an outbreak (i.e. at different times the number of cases, hospitalisations and/or deaths may be attributed different importance); consideration of these behavioural feedbacks would require refinements to our model to include severe case outcomes (such as hospitalisations and deaths).
@@
Associated with our idealised assumptions on testing capacity, another direction for enhancing the model is onboarding additional data steams for parameter estimation, such as for the test-seeking propensity. As an initial entry point, publicly available data at a coarse demographic-level may be provided by public health agencies. Using the COVID-19 pandemic as an example setting, in the UK statistics for testing were made available in weekly reports [36]. Yet, such data may not have the sufficient detail to capture the multiple potential influences on testing uptake, including employment conditions, geographical location, population density, beliefs about testing, health literacy, region, index of multiple deprivation (IMD) or other indicator of deprivation, age and/or gender. Pseudonymised individual-level data would be very beneficial for capturing such signals, which may then be used in data-driven models. Such individual-level data would likely not be publicly available for data protection reasons, but instead be held in secure data environments; for example, pseudonymised individual-level SARS-CoV-2 lateral flow testing data for Liverpool in the UK was used by Green et al. [37] to evaluate social and spatial inequalities.
@@ Acknowledgements
EMH is affiliated to the NIHR Health Protection Research Unit in Emerging and Zoonotic Infections (NIHR207393). The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care. EMH is also funded by The Pandemic Institute, formed of seven founding partners: The University of Liverpool, Liverpool School of Tropical Medicine, Liverpool John Moores University, Liverpool City Council, Liverpool City Region Combined Authority, Liverpool University Hospital Foundation Trust, and Knowledge Quarter Liverpool (EMH is based at The University of Liverpool). The views expressed are those of the author(s) and not necessarily those of The Pandemic Institute. TH is supported by the Wellcome Trust (Grant number 227438/Z/23/Z). MGR is supported by the Marsden Fund under contract 20-MAU-50. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme “Modelling and inference for pandemic preparedness”, where work on this paper was initiated. This work was supported by EPSRC grant EP/Z000580/1. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@
Code availability statement
The code associated with this study is available at:
https://github.com/Matthew-Ryan1995/BaD_testing_and_isolation.
Appendix A Stability of the infection-free equilibrium
We will first derive the Jacobian of the system and evaluate its eigenvalues at the infection-free equilibrium, thereby confirming Proposition 3.1 that stability holds when $\mathcal{R}_{0}<1$.
For notational convenience, we define
and analogously for behavers:
with
In this notation, the BaD model can be written in matrix form as
where $\mathbf{u}_{1}$ is the first standard basis vector, having a first element equal to $1$ and all other elements equal to 0.
Next, the Jacobian $\mathbf{J}$ of the full system (36) has the block form
where $\mathbf{J}^{0}$ is a $2\times 2$ matrix given by
$\mathbf{J}^{n0}$ and $\mathbf{J}^{b0}$ are $4\times 2$ matrices given by
$\mathbf{J}^{n0}$ $=\left(\begin{array}[]{cc}\dfrac{\partial\omega}{\partial B}S_{N}+\dfrac{\partial\alpha}{\partial N}S_{B}+\nu&-q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}S_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}E_{N}+\dfrac{\partial\alpha}{\partial N}E_{B}&q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}A_{N}+\dfrac{\partial\alpha}{\partial N}A_{B}&-\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}I_{N}+\dfrac{\partial\alpha}{\partial N}I_{B}&-\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right)$
$\mathbf{J}^{b0}$ $=\left(\begin{array}[]{cc}-\dfrac{\partial\omega}{\partial B}S_{N}-\dfrac{\partial\alpha}{\partial N}S_{B}-\nu&-q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}S_{N}-\nu\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}E_{N}-\dfrac{\partial\alpha}{\partial N}E_{B}&q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}A_{N}-\dfrac{\partial\alpha}{\partial N}A_{B}&\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}I_{N}-\dfrac{\partial\alpha}{\partial N}I_{B}&\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right)$
and the other four sub-matrices $\mathbf{J}^{nn}$, $\mathbf{J}^{nb}$, $\mathbf{J}^{bn}$ and $\mathbf{J}^{bb}$ have dimension $4\times 4$, and are given by
At the infection-free steady state, we have $S_{N}=N_{0}$, $S_{B}=B_{0}$, with $B_{0}=1-N_{0}.$ It follows that
$\omega(B_{0},0)=\omega_{1}B_{0}+\omega_{3},\quad\alpha(N_{0})=\alpha_{1}N_{0}+\alpha_{2}.$
We also have
$\dfrac{\partial\omega}{\partial B}=\omega_{1},\quad\dfrac{\partial\omega}{\partial T}=\omega_{2},\quad\dfrac{\partial\alpha}{\partial N}=\alpha_{1}.$
Substituting into the Jacobian blocks gives, we obtain
$\mathbf{J}^{n0}$ $=\left(\begin{array}[]{cc}\omega_{1}N_{0}+\alpha_{1}B_{0}+\nu&-q_{T}\beta N_{0}-\omega_{2}N_{0}\\ 0&q_{T}\beta N_{0}\\ 0&0\\ 0&0\end{array}\right)$
$\mathbf{J}^{b0}$ $=\left(\begin{array}[]{cc}-\omega_{1}N_{0}-\alpha_{1}B_{0}-\nu&-q_{B}q_{T}\beta B_{0}+\omega_{2}N_{0}-\nu\\ 0&q_{B}q_{T}\beta B_{0}\\ 0&0\\ 0&0\end{array}\right)$
$\text{and}\quad\mathbf{J}^{bb}$ $=\left(\begin{array}[]{cccc}-\alpha(N_{0})-\nu&-\nu&-q_{A}q_{B}\beta B_{0}-\nu&-q_{B}\beta B_{0}-\nu\\ 0&-\sigma-\alpha(N_{0})&q_{A}q_{B}\beta B_{0}&q_{B}\beta B_{0}\\ 0&p_{A}\sigma&-\gamma-\alpha(N_{0})&0\\ 0&\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma&0&-\gamma-\alpha(N_{0})\end{array}\right).$
The infection-free steady state where $S_{N}=N_{0}$ and $S_{B}=B_{0}$ follows standard results for disease dynamics without a behavioural component and so is stable when $\mathcal{R}_{0}<1$.
Appendix B Derivation of the endemic steady states
Here we present details on derivations underlying Proposition 3.2. When infection is present and ignoring higher order terms, the dynamics of behaviour are governed by
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,T)N+\alpha(N)(B-T)$
$=(\omega_{1}-\alpha_{1})N^{2}-(\omega_{2}T+\omega_{1}+\omega_{3}+\alpha_{2}-\alpha_{1}(1-T))N+\alpha_{2}(1-T).$ (37)
At equilibrium, $N^{*}$ satisfies $G(N^{*})=0$ (Equation (30)). Since $G(0)=\alpha_{2}(1-T^{*})>0$ and $G(1)=-[(\omega_{2}+\alpha_{1}+\alpha_{2})T^{*}+\omega_{3}]<0$, there exists a unique solution $N^{*}\in(0,1)$ depending on $T^{*}$. The behaviour prevalence is then $B^{*}=1-N^{*}$.
For $X\in\{S,E,A,I,R\}$ write $X=X_{N}+X_{B}$ and let the total symptomatic be $O=I+T$. Summing Equations (6), (11) and (12) together gives
$\frac{\mbox{d}O}{\mbox{d}t}=(1-p_{A})\sigma(E_{N}+E_{B})-\gamma O.$
Similarly, adding Equations (5) and (10), Equations (7) and (13) gives
$\frac{\mbox{d}A}{\mbox{d}t}$ $=p_{A}\sigma E-\gamma A$
$\frac{\mbox{d}R}{\mbox{d}t}$ $=\gamma(I+A+T)-\nu R.$
Hence, fixing $O^{*}=I^{*}+T^{*}$, setting the equations above to zero yields
$E^{*}$ $=E_{N}^{*}+E_{B}^{*}=\frac{\gamma}{(1-p_{A})\sigma}O^{*}$ (38)
$A^{*}$ $=A_{N}^{*}+A_{B}^{*}=\frac{p_{A}}{(1-p_{A})}O^{*}$ (39)
$R^{*}$ $=R_{N}^{*}+R_{B}^{*}=\frac{\gamma}{\nu(1-p_{A})}O^{*}$ (40)
$S^{*}$ $=S_{N}^{*}+S_{B}^{*}=1-(E^{*}+A^{*}+O^{*}+R^{*}).$ (41)
Further, knowing $T^{*}$ (and hence $N^{*}$ and $B^{*}=1-N^{*}$), the equilibrium rates $\lambda^{*},\omega^{*}$ and $\alpha^{*}$ are in Equations (31)-(33).
Next, we derive the behavioural steady states $S_{B}^{*},E_{B}^{*},A_{B}^{*},I_{B}^{*}$ and $R_{B}^{*}$.
Setting Equations (4) and (9) to zero gives
Adding Equations (42) and (43) gives
$E_{B}^{*}=\frac{q_{B}\lambda^{*}S^{*}+(\omega^{*}-q_{B}(\omega^{*}+\sigma))E^{*}}{(1-q_{B})(\alpha^{*}+\omega^{*}+\sigma)}.$
Setting Equations (10) and (11) to zero respectively gives
$A_{B}^{*}$ $=\frac{p_{A}\sigma E_{B}^{*}+\omega^{*}A^{*}}{\gamma+\alpha^{*}+\omega^{*}}$
$I_{B}^{*}$ $=\frac{(1-p_{A})(1-p_{T})\sigma E_{B}^{*}+\omega^{*}I^{*}}{\gamma+\alpha^{*}+\omega^{*}}.$
Further, setting Equation (13) to zero gives
$R_{B}^{*}=\frac{\gamma(A_{B}^{*}+I_{B}^{*}+T^{*})+\omega^{*}R^{*}}{\alpha^{*}+\omega^{*}+\nu}\,.$
It follows that
$S_{B}^{*}=B^{*}-(E_{B}^{*}+A_{B}^{*}+I_{B}^{*}+T^{*}+R_{B}^{*}).$
Given $T^{*}$ and $O^{*}$, the non-behaver states $S_{N}^{*},E_{N}^{*},A_{N}^{*},I_{N}^{*}$ and $R_{N}^{*}$ are obtained by subtraction:
$S_{N}^{*}$ $=S^{*}-S_{B}^{*}$
$E_{N}^{*}$ $=E^{*}-E_{B}^{*}$
$A_{N}^{*}$ $=A^{*}-A_{B}^{*}$
$R_{N}^{*}$ $=R^{*}-R_{B}^{*}$
$I_{N}^{*}$ $=N^{*}-(S_{N}^{*}+E_{N}^{*}+A_{N}^{*}+R_{N}^{*}).$
Finally, setting Equations (3) and (12) to zero gives the simultaneous Equations (34) and (35) which fixes $I^{*}$ and $T^{*}$ (and hence $O^{*}$).
Supplementary Figures
Supplementary Figure 8: Change in (a) final size and (b) peaks for different $\mathcal{R}_{0}$} values. The dashed, green line represented the observed epidemic ($T$), the dotted, orange line represents the symptomatic epidemic ($O=I+T$), and the solid, red line represents the total epidemic ($O+A$).
Supplementary Figure 9: Transient dynamics of the behaviour and disease model for testing and isolation with baseline parameters from Table 1 in the main text. (a) The transient dynamics for each epidemiological compartment. (b) The transient dynamics for each of the five infectious classes, where dotted lines present behaviours. (c) The transient dynamics of behaviours ($B(t)$). (d) A comparison of the symptomatic infection prevalence $O(t)$ across: 1) the baseline parameters with dynamic behaviour, 2) A model with no behaviour at all, and 3) a model where behaviour is held static throughout the whole simulation.
Supplementary Figure 10: Impact of test effectiveness ($p_{T}$), isolation effectiveness ($1-q_{T}$), initial condition of behaviour (B(0){B(0)}) and the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$}) on observed symptomatic (T{T}), undetected symptomatic (I{I}), and total symptomatic (O=I+T{O=I+T}) peaks. The colour bars represent the peak infection prevalence as a percentage of the population and are consistent across each row. Note, we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ due to the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey lines show $\mathcal{R}_{0}=1$ and the black cross indicates the baseline parameter values from Table 1 in the manuscript.
Supplementary Figure 11: Phase diagrams for different behavioural parameters fixing $\mathcal{R}_{0}^{D}=3.28$} in the endemic regime. Column one corresponds to changes in social influence ($\omega_{1}$), column two corresponds to changes in perception of illness threat ($\omega_{2}$), and column three corresponds to spontaneous uptake ($\omega_{3}$), respectively. The rows show the susceptible versus observed infection ($S$ vs $T$), the susceptible versus true infection ($S$ vs $O+A$), and the behaviour versus observed infection ($B$ vs $T$) phase planes. The dashed line represents near removal of the behaviour construct, the dotted lines represent a reduction of the base line value, the solid line represents the baseline values, and the dash-dot lines represent an increase in the baseline values. The solid lines all represent the baseline values from Table 1 in the manuscript, ensuring the solid lines are comparable across each column. Each simulation was run with the initial conditions $S_{B}(0)=10^{-6},I_{N}(0)=10^{-6},S_{N}=1-S_{B}-I_{N}$ and all other compartments empty: this ensured the early-stage dynamics of each model were approximately comparable.